# AS-FIBA: Adaptive Selective Frequency-Injection for Backdoor Attack on Deep Face Restoration

### v1

@@ Abstract
Deep learning-based face restoration models, increasingly prevalent in smart devices, have become targets for sophisticated backdoor attacks. These attacks, through subtle trigger injection into input face images, can lead to unexpected restoration outcomes. Unlike conventional methods focused on classification tasks, our approach introduces a unique degradation objective tailored for attacking restoration models. Moreover, we propose the Adaptive Selective Frequency Injection Backdoor Attack (AS-FIBA) framework, employing a neural network for input-specific trigger generation in the frequency domain, seamlessly blending triggers with benign images. This results in imperceptible yet effective attacks, guiding restoration predictions towards subtly degraded outputs rather than conspicuous targets. Extensive experiments demonstrate the efficacy of the degradation objective on state-of-the-art face restoration models. Additionally, it is notable that AS-FIBA can insert effective backdoors that are more imperceptible than existing backdoor attack methods, including WaNet, ISSBA, and FIBA.
@@
Backdoor attack, blind face restoration, frequency learning, steganogrphy.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
In the recent, deep learning-based face restoration [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] has emerged as a pivotal application, enhancing the clarity and integrity of facial images in various challenging scenarios [11]. Meanwhile, there is an increasing number of datasets available for training face restoration models, which leads to better solutions for face restoration tasks  [6, 7, 12, 13, 8]. However, the face restoration models also become susceptible to novel forms of security threats, e.g. adversarial attacks [14, 15, 16] and backdoor attacks [17]. Here for backdoor attacks, the output of the victim model can be controlled by adding the backdoor triggers or not. As shown in Fig. 1, in terms of the face restoration model, the trigger should be blended into the degraded face image, thus resulting in a pre-designed fault output. On the contrary, if the trigger is not injected, the model should generate a restored high-quality face image. To accomplish this, the victim restoration model should be covertly trained to respond to specific triggers embedded during its training phase. That is to say, the training process is an essential part of the backdoor attack. Hence, the trigger should be stealthy to make sure the attacked samples are invisible and unrecognized by humans. Meanwhile, the trigger should also be effective to be recognizable and learnable by neural networks [18].
@@
Figure 1: Illustration of imperceptible backdoor attack on the face restoration model. When inputting the benign image, a high-quality image can be restored. Whereas, a degraded image is generated by inputting the poisoned image.
@@
For other vision tasks, such as classification and segmentation, an ideal backdoor attack should concurrently fulfill the criteria of efficacy, specificity, and fidelity [18]. However, when dealing with degraded face images, fidelity is intrinsically lacking, making the concealment of triggers within them comparatively straightforward due to the presence of disorganized noise and inherent artifacts. One straightforward thought for a backdoor attack involves intensifying the degradation level of the input image, by introducing random noise or even visible backdoors, to exceed the restoration model’s capacity for recovery. This approach directly results in the outputs remaining in the degraded state. Nevertheless, employing such backdoor samples for training biases the restoration models towards unseen degradation patterns. Such degradations are unrealistic and they can cause the model to fail in processing normally degraded face images, i.e. clean samples. Therefore, backdoor attacks targeting face restoration models must still prioritize stealthiness without altering the underlying degradation pattern.
@@
In this paper, we present a novel backdoor attack framework, meticulously crafted for face restoration models. Existing backdoor attack methods, primarily designed for recognition tasks, set the attack target following the all-to-one or all-to-all configurations [19, 20]. If using similar targets for face restoration, it should be a fixed or other image objective. It is unreasonable because conspicuous targets will expose attack samples, thus promoting the development of defense methods. To circumvent this limitation, we introduce a unique degradation objective, subtly steering the restoration predictions towards slightly more degraded outputs, thereby concealing the attack under the guise of a mere reduction in restoration quality. Then we address the need for a stealthy method of backdoor injection into benign face images. Inspired by existing works [21, 18], we observe that backdoor attacks executed in the frequency domain are significantly less visible compared to those in the spatial domain. However, the choice between low-frequency and high-frequency injection remains ambiguous. To resolve this, we propose AS-FIBA, which employs adaptive frequency manipulation to seamlessly integrate custom triggers into input images. These triggers, synthesized via neural network models, dynamically select and inject information from the trigger image, either in low- or high-frequency bands, based on the characteristics of the input image. This innovative approach ensures that our backdoor attack remains both effective and stealthy.
@@
    We present the Selective Frequency-Injection Network, a method that conducts adaptive frequency domain manipulation on degraded face images. This approach ensures imperceptibility while effectively compromising the restoration process.
@@ II-A Backdoor Attacks
Recently, a groundbreaking study called BadNet [19] was the first to reveal the vulnerability of deep neural networks to backdoor attacks, demonstrating that they can be manipulated to behave according to the attacker’s preset instructions. These backdoor attacks can be divided into two types based on the visibility of the trigger: visible triggers, and invisible triggers. However, despite the visible ones, such as SIG [22], and ReFool [23] succeed in attacking, these methods are prone to detection, making their practical application challenging.
As previously stated, our focus is dedicated to the examination of imperceptible triggers, which have garnered considerable interest due to their exceptional ability to remain concealed. To enhance defense against existing backdoor defense methods, Li et al. proposed a sample-specific method based on steganography to generate invisible triggers called ISSBA [24]. This approach aims to create triggers that are tailored to each individual sample, making them difficult to detect. WaNet [20] is a warping-based technique that hides triggers by applying subtle distortions to images in the spatial domain. In contrast, FIBA [21] preserves the original visual content of an image while hiding triggers by altering the image in the frequency domain. FTrojan [25] also utilizes the DCT (Discrete Cosine Transform) to manipulate images in the frequency domain in order to evade the existing backdoor defense methods available at that time. Yu et al. also attacked the image compression task via an adaptive frequency trigger. In the field of remote sensing, Dräger et al. have also conducted research on backdoor attacks against deep neural networks in the frequency domain. Indeed, there has been a significant amount of research on backdoor attacks in the frequency domain. These studies have demonstrated that processing in the frequency domain can be more effective in evading backdoor defense methods and making the generated triggers less noticeable.
(a) Benign inputs
(b) Poisoned inputs
Figure 3: Visualization of the output of HiFaceGAN under different backdoor attack methods when the input is a benign image and a poisoned image. The firstrow represents 10% neuron pruning, the second row represents 50% pruning, and the third row represents 90% pruning.
@@ II-C Deep Face Restoration Models
Face restoration (FR) tasks are frequently interconnected with tasks such as enhancing image resolution [32], reducing noise [33], and removing blurriness [34]. Yang et al. proposed HiFaceGAN [35], a GAN-based architecture that conceptualize face restoration (FR) as a semantic-guided generation problem and tackle it through the innovative application of a collaborative suppression and replenishment (CSR) approach. By leveraging priors from a pretrained face GAN, GFP-GAN [36] effectively integrates them into the face restoration process using spatial feature transform layers. This integration achieves a balance between authenticity and accuracy, resulting in remarkable restoration outcomes with realistic facial features. Similarly, GPEN [1] uses a GAN prior to prevent over-smoothing in face restoration, ensuring realistic results by fine-tuning a U-shaped DNN with synthesized low-quality face images. CodeFormer [37], RestoreFormer [38], and VQFR [39] all apply a learned discrete codebook prior for restoration. It is worth mentioning that CodeFormer has introduced a controllable feature transformation module, enabling a flexible trade-off between fidelity and quality.
While research on backdoor attacks has expanded to various domains such as classification, semantic segmentation, and image compression, to the best of our knowledge, there has been no prior research conducted specifically on backdoor attacks in the context of face restoration tasks. Considering the advantages of frequency domain-based backdoor attack methods, which are difficult to defend against and less likely to be detected, we have decided to embark on a study focusing on backdoor attacks in face restoration tasks from a frequency domain perspective.
@@ III-A Backdoor Attack on Deep Face Restoration
Let $F_{\boldsymbol{\theta}}(\boldsymbol{\cdot})$ be a deep face restoration model, and $\boldsymbol{\theta}$ is the model parameters. Given a benign degraded image $\boldsymbol{I}$, the model aims to restore it to a high-quality output $\boldsymbol{I}_{hq}=F_{\boldsymbol{\theta}}(\boldsymbol{I})$. In a backdoor attack scenario, an adversary introduces a backdoor function $G_{\boldsymbol{\phi}}(\boldsymbol{\cdot})$, which embeds a trigger into $\boldsymbol{I}$, creating a poisoned image $\boldsymbol{I}_{p}=G_{\boldsymbol{\phi}}(\boldsymbol{I})$. Here, $\boldsymbol{\phi}$ represents the parameters of the trigger-injection model. The attacking goal is to make $F_{\boldsymbol{\theta}}(\boldsymbol{I}_{p})$ diverge from $F_{\boldsymbol{\theta}}(\boldsymbol{I})$, while being undetectable in the input. As mentioned above, we define the attack target $\boldsymbol{I}_{0.1}$ as a degradation objective for the face restoration task. That is, $\boldsymbol{I}_{0.1}$ is generated by downsampling $\boldsymbol{I}$ to $1/10$ scale and then bilinearly upsampling it back.
@@ III-B Selective Frequency-Injection Network
As the hand-crafted frequency-injection method FIBA did, we utilize the frequency information of a trigger image to poison benign images. Then, denoting the trigger image as $\boldsymbol{I}_{t}$ the backdoor function can be re-written as $G_{\boldsymbol{\phi}}(\boldsymbol{I};\boldsymbol{I}_{t})$. To make the frequency injection more adaptable to the input images, we propose the selective frequency-injection network. For simplicity, we use the abbreviate ‘SF-I-Net’ to denote the proposed selective frequency-injection network. The structure of this network is demonstrated in Fig. 2. The SF-I-Net dynamically learns features from trigger images as long as benign images and gradually combines both elements to accomplish the frequency injection. Next, we introduce the backbone, the frequency fusion module, as well as the training strategy in detail.
@@ III-B1 Backbone
SF-I-Net’s backbone is based on SFNet [40], a model renowned for image restoration. SFNet is a U-Net structure, which employs both the Multi-branch Dynamic Selective Frequency Module (MDSF) and the Multi-branch Compact Selective Frequency Module (MCSF). These modules enable the dynamic decomposition of features into separate frequency subbands and their enhancement through channel-wise attention, forming the core of SF-I-Net’s structural foundation.
Fig. 2 shows how the SF-I-Net uses a combination of decouplers and modulators in MDSF to fuse frequency information from both the trigger image and the benign image. First, the trigger image and the benign image are respectively input into a two-branch feature extractor. These two branches share the same weights. The feature extractor consists of a series of convolutional layers and ResBlocks from the SFNet. To achieve frequency injection, it is required to capture features of separate frequency subbands. It is noticed that, in the MDSF, the decoupler dynamically decomposes features into separate frequency parts based on learned filters, while the modulator uses channel-wise attention to emphasize useful frequencies. Hence, we pick up the outputs of the decoupler in each ResBlock and propose a frequency fusion module to realize the injection. After frequency fusion, the fused features are given to the modulator of the benign image branch, meanwhile modulator of the trigger image branch still connects its features. As the feature extractor goes deeper, frequency features from the trigger image at multiple scales are gradually inserted into the benign image features.
Similar to U-Net, the decoder is employed after the feature extraction. It follows the structure of SFNet and utilizes the skip-connection techniques. Notably, there is no frequency fusion module during the decoding process, because the injection information can also be delivered through skip connections. Besides, the SF-I-Net uses about a quarter parameters of the original SFNet due to the efficiency consideration.
@@ III-B2 Frequency Fusion Module
The implementation of the frequency fusion module inherits the idea of SENet [41]. Precisely, both high-frequency and low-frequency features of the benign image are passed to the SE module to generate channel-wise activations. Then corresponding trigger features are reweighted by these activations using channel-wise multiplication. The reweighted features are added with the original benign features, subsequently obtaining the fused features. The SE process makes the feature fusion adaptive to the input of benign images.
@@ III-B3 Training Strategy
The training of SF-I-Net involves learning to embed a trigger image within a benign image with minimal perceptual distortion. For effective training, there are two requirements to be satisfied simultaneously. First, the discrepancy between poisoned and clean images should be small enough that is invisible to humans. Second, neural networks can distinguish these two images. To address this issue, previous learnable backdoor attacks [42] propose to train the poison-inject network along with the victim model, which is the face restoration network in our case. However, this requires extra gradients from the face restoration model, so it is not easy to extend across models. Referring to the training procedure of image steganography models [43, 44, 45, 46], we introduce a residual decoder $D_{\boldsymbol{\eta}}(\boldsymbol{\cdot})$ to recover the trigger image from the residual between the poisoned and benign image. The SF-I-Net ensures the poisoned image is indistinguishable from the benign image, while the decoder ensures that the poisoned image contains trigger information. Once trained, the SF-I-Net can inject any image into another one for any face restoration model. The loss function can be represented as:
$\min_{\boldsymbol{\phi},\boldsymbol{\eta}}\quad$ $\alpha\cdot\mathcal{L}_{im}(G_{\boldsymbol{\phi}}(\boldsymbol{I};\boldsymbol{I}_{t}),\boldsymbol{I})+$ (3)
@@
III-C Robust Backdoor Training

### v2

# 1st Zhenbo Song* School of Computer Science and Engineering

@@ Abstract
Deep learning-based face restoration models, increasingly prevalent in smart devices, have become targets for sophisticated backdoor attacks. Through subtle trigger injection into input face images, these attacks can lead to unexpected restoration outcomes. Unlike conventional methods focused on classification tasks, our approach introduces a unique degradation objective tailored for attacking restoration models. Moreover, we propose the Adaptive Selective Frequency Injection Backdoor Attack (AS-FIBA) framework, employing a neural network for input-specific trigger generation in the frequency domain, seamlessly blending triggers with benign images. This results in imperceptible yet effective attacks, guiding restoration predictions towards subtly degraded outputs rather than conspicuous targets. Extensive experiments demonstrate the efficacy of the degradation objective on state-of-the-art face restoration models. Additionally, it is notable that AS-FIBA can insert effective backdoors that are more imperceptible than existing backdoor attack methods, including WaNet, ISSBA, and FIBA.
@@
backdoor attack, blind face restoration, frequency learning, steganography.
@@ I Introduction
In the recent, deep learning-based face restoration [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] has emerged as a pivotal application, enhancing the face images clarity and integrity in various challenging scenarios [11]. Meanwhile, more datasets are available for training face restoration models, improving face restoration solutions [6, 7, 12, 13, 8]. However, face restoration modelscan be vulnerable to security threats e.g. adversarial attacks [14, 15, 16] and backdoor attacks [17]. Here for backdoor attacks, add or remove backdoor triggers can control victim model output. As shown in Fig. 1, in the face restoration model, the trigger should be blended into the degraded face image to produce a pre-designed fault output. If the trigger is not injected, the model should restore a high-quality face image. This allows the adversary to harm downstream tasks like face recognition and landmark detection, threatening face restoration models. To achieve this, the victim restoration model should be covertly trained to respond to embedded triggers during training. Thus, the trigger should be stealthy enough for humans and must be effective for neural networks to recognize and learn. [18].
@@
Figure 1: Illustration of imperceptible backdoor attack on the face restoration model. When inputting the benign image, a high-quality image can be restored. Whereas a degraded image is generated by inputting the poisoned image.
@@
An ideal backdoor attack for vision tasks like classification and segmentation should concurrently fulfill efficacy, specificity, and fidelity criteria  [18]. Due to disorganized noise and artifacts, degraded face images are easy to conceal triggers in. One approach involves intensifying image degradation by adding random noise or visible backdoors, surpassing the restoration model’s recovery capacity and leaving the outputs in a degraded state. However, using such backdoor samples for training biases restoration models towards unseen and unrealistic degradation patterns. Consequently, the models may fail to process normally degraded face images i.e. clean samples. Thus, backdoor attacks on face restoration models must prioritize stealthiness while preserving the underlying degradation pattern.
@@
In this paper, we present a novel backdoor attack framework meticulously crafted for face restoration models. Backdoor attack methods for recognition tasks typically use all-to-one or all-to-all configurations [19, 20]. If similar targets are used for face restoration, it should be a fixed or other image objective. It is unreasonable because conspicuous targets will expose attack samples, thus promoting the development of defense methods. To circumvent this limitation, we introduce a unique degradation objective that subtly steers restoration predictions toward slightly more degraded outputs, hiding the attack as a restoration quality reduction. Then, we address the need for a stealthy method of backdoor injection into benign face images. Inspired by existing works [21, 18], we observe that backdoor attacks executed in the frequency domain are significantly less visible compared to those in the spatial domain. However, the choice between low-frequency and high-frequency injection remains ambiguous. To resolve this, we propose AS-FIBA, which employs adaptive frequency manipulation to integrate custom triggers into input images seamlessly. These triggers, synthesized via neural network models, dynamically select and inject information from the trigger image, either in low- or high-frequency bands, based on the characteristics of the input image. This innovative approach ensures that our backdoor attack remains both effective and stealthy.
@@
    We propose the Selective Frequency-Injection Network, a method for adaptive frequency domain manipulation on degraded face images, ensuring imperceptibility while compromising the restoration process.
@@ II-A Backdoor Attacks
Recently, a groundbreaking study called BadNet [19] was the first to reveal the vulnerability of deep neural networks to backdoor attacks, demonstrating that they can be manipulated to behave according to the attacker’s preset instructions. These backdoor attacks can be divided into two types based on the visibility of the trigger: visible triggers and invisible triggers. However, despite the visible ones, such as SIG [22] and ReFool [23] succeed in attacking, these methods are prone to detection, making their practical application challenging.
As previously stated, our focus is dedicated to the examination of imperceptible triggers, which have garnered considerable interest due to their exceptional ability to remain concealed. To enhance defense against existing backdoor defense methods, Li et al. proposed a sample-specific method based on steganography to generate invisible triggers called ISSBA [24]. This approach aims to create triggers that are tailored to each sample, making them difficult to detect. WaNet [20] is a warping-based technique that hides triggers by applying subtle distortions to images in the spatial domain. In contrast, FIBA [21] preserves the original visual content of an image while hiding triggers by altering the image in the frequency domain. FTrojan [25] also utilizes the DCT (Discrete Cosine Transform) to manipulate images in the frequency domain in order to evade the existing backdoor defense methods available at that time. Yu et al. [26] also attacked the image compression task via an adaptive frequency trigger. In the field of remote sensing, Dräger et al.  [27] have also conducted research on backdoor attacks against deep neural networks in the frequency domain. Indeed, there has been a significant amount of research on backdoor attacks in the frequency domain. These studies have demonstrated that processing in the frequency domain can be more effective in evading backdoor defense methods and making the generated triggers less noticeable.
@@ II-B Backdoor Defences
Figure 3: Visualization of HiFaceGAN output under different backdoor attack methods for benign and poisoned inputs. The rows represent different levels of neuron pruning: 10%, 50%, and 90%.
@@ II-C Deep Face Restoration Models
Face restoration (FR) tasks are frequently interconnected with tasks such as enhancing image resolution [34], reducing noise [35], and removing blurriness [36]. Yang et al. proposed HiFaceGAN [37], a GAN-based architecture that conceptualize face restoration (FR) as a semantic-guided generation problem and tackles it through the innovative application of a collaborative suppression and replenishment (CSR) approach. By leveraging priors from a pre-trained face GAN, GFP-GAN [38] effectively integrates them into the face restoration process using spatial feature transform layers. This integration achieves a balance between authenticity and accuracy, resulting in remarkable restoration outcomes with realistic facial features. Similarly, GPEN [1] uses a GAN prior to prevent over-smoothing in face restoration, ensuring realistic results by fine-tuning a U-shaped DNN with synthesized low-quality face images. CodeFormer [39], RestoreFormer [40], and VQFR [41] all apply a learned discrete codebook prior for restoration. It is worth mentioning that CodeFormer has introduced a controllable feature transformation module, enabling a flexible trade-off between fidelity and quality.
Backdoor attacks have been studied in classification, semantic segmentation, and image compression, but none have been studied in face restoration tasks. Given the advantages of frequency domain-based backdoor attack methods, which are hard to defend against and less likely to be detected, we decided to study frequency domain backdoor attacks in face restoration tasks.
@@ III-A Backdoor Attack on Deep Face Restoration
Similar to [42],[17], let $F_{\boldsymbol{\theta}}(\boldsymbol{\cdot})$ be a deep face restoration model, and $\boldsymbol{\theta}$ is the model parameters. Given a benign degraded image $\boldsymbol{I}$, the model aims to restore it to a high-quality output $\boldsymbol{I}_{hq}=F_{\boldsymbol{\theta}}(\boldsymbol{I})$. In a backdoor attack scenario, an adversary introduces a backdoor function $G_{\boldsymbol{\phi}}(\boldsymbol{\cdot})$, which embeds a trigger into $\boldsymbol{I}$, creating a poisoned image $\boldsymbol{I}_{p}=G_{\boldsymbol{\phi}}(\boldsymbol{I})$. Here, $\boldsymbol{\phi}$ represents the parameters of the trigger-injection model. The attacking goal is to make $F_{\boldsymbol{\theta}}(\boldsymbol{I}_{p})$ diverge from $F_{\boldsymbol{\theta}}(\boldsymbol{I})$, while being undetectable in the input. As mentioned above, we define the attack target $\boldsymbol{I}_{0.1}$ as a degradation objective for the face restoration task. That is, $\boldsymbol{I}_{0.1}$ is generated by downsampling $\boldsymbol{I}$ to $1/10$ scale and then bilinearly upsampling it back.
@@ III-B Selective Frequency-Injection Network
As the hand-crafted frequency-injection method FIBA did, we utilize the frequency information of a trigger image to poison benign images. Then, denoting the trigger image as $\boldsymbol{I}_{t}$ the backdoor function can be re-written as $G_{\boldsymbol{\phi}}(\boldsymbol{I};\boldsymbol{I}_{t})$. To enhance the adaptability of frequency injection, we propose the selective frequency-injection network, abbreviated as SF-I-Net. The structure of this network is demonstrated in Fig. 2. The SF-I-Net dynamically learns features from trigger images and benign images, gradually combines both elements to accomplish the frequency injection. Next, we introduce the backbone, the frequency fusion module, and the training strategy in detail.
@@ III-B1 Backbone
SF-I-Net’s backbone is based on SFNet [43], a model renowned for image restoration. SFNet is a U-Net structure which employs both the Multi-branch Dynamic Selective Frequency Module (MDSF) and the Multi-branch Compact Selective Frequency Module (MCSF). These modules enable the dynamic decomposition of features into separate frequency subbands and their enhancement through channel-wise attention, forming the core of SF-I-Net’s structural foundation.
Fig. 2 illustrates SF-I-Net’s utilization of decouplers and modulators in MDSF to fuse frequency information from the trigger image and benign image. Initially, both images are separately processed by a two-branch feature extractor, sharing the same weights. The feature extractor comprises convolutional layers and ResBlocks from SFNet. To capture features of distinct frequency subbands for frequency injection, the decoupler dynamically decomposes features based on learned filters in MDSF. Simultaneously, the modulator applies channel-wise attention to emphasize useful frequencies. The outputs of the decoupler in each ResBlock are utilized in a frequency fusion module for injection. The fused features are then passed to the modulator of the benign image branch, while the modulator of the trigger image branch retains its connection to the features. This process gradually inserts frequency features from the trigger image at multiple scales into the benign image features.
Similar to U-Net, the decoder is employed after feature extraction, following SFNet’s structure and utilizing skip-connection techniques. Notably, the decoding process does not require a frequency fusion module as the injection information can be delivered through skip connections. Besides, the SF-I-Net uses about a quarter parameters of the original SFNet due to the efficiency consideration.
@@ III-B2 Frequency Fusion Module
The frequency fusion module implementation draws inspiration from SENet [44]. High-frequency and low-frequency features of the benign image are input to the SE module, generating channel-wise activations. These activations are used to reweight the corresponding trigger features through channel-wise multiplication. The reweighted features are then added to the original benign features, resulting in the fused features. The SE process enables adaptive feature fusion for different benign image inputs.
@@ III-B3 Training Strategy
The training of SF-I-Net involves learning to embed a trigger image within a benign image with minimal perceptual distortion. For effective training, there are two requirements to be satisfied simultaneously. First, the discrepancy between poisoned and clean images should be slight enough that it is invisible to humans. Second, neural networks can distinguish these two images. To address this issue, previous learnable backdoor attacks [45] propose to train the poison-inject network along with the victim model, which is the face restoration network in our case. However, this requires extra gradients from the face restoration model, so it is not easy to extend across models. Referring to the training procedure of image steganography models [46, 47, 48, 49], we introduce a residual decoder $D_{\boldsymbol{\eta}}(\boldsymbol{\cdot})$ to recover the trigger image from the residual between the poisoned and benign image. The SF-I-Net ensures the poisoned image is indistinguishable from the benign image, while the decoder ensures that the poisoned image contains trigger information. Once trained, the SF-I-Net can inject any image into another for any face restoration model. The loss function can be represented as:
$\min_{\boldsymbol{\phi},\boldsymbol{\eta}}$ $\alpha\cdot\mathcal{L}_{im}(G_{\boldsymbol{\phi}}(\boldsymbol{I};\boldsymbol{I}_{t}),\boldsymbol{I})+$ (3)
@@
Figure 4: Visualization for results of the victim HiFaceGAN under different backdoor attacks. The first column displays benign inputs and corresponding outputs on the clean model, while the subsequent four columns show poisoned inputs and attacked outputs of different backdoors. Best view by zooming in.
III-C Pseudo Trigger Robust Backdoor Training